# UPose3D: Uncertainty-Aware 3D Human Pose Estimation with Cross-View and Temporal Cues

### v2

@@ Abstract
We introduce UPose3D, a novel approach for multi-view 3D human pose estimation, addressing challenges in accuracy and scalability. Our method advances existing pose estimation frameworks by improving robustness and flexibility without requiring direct 3D annotations. At the core of our method, a pose compiler module refines predictions from a 2D keypoints estimator that operates on a single image by leveraging temporal and cross-view information. Our novel cross-view fusion strategy is scalable to any number of cameras, while our synthetic data generation strategy ensures generalization across diverse actors, scenes, and viewpoints. Finally, UPose3D leverages the prediction uncertainty of both the 2D keypoint estimator and the pose compiler module. This provides robustness to outliers and noisy data, resulting in state-of-the-art performance in out-of-distribution settings. In addition, for in-distribution settings, UPose3D yields a performance rivaling methods that rely on 3D annotated data, while being the state-of-the-art among methods relying only on 2D supervision.
Index Terms:
Markerless Motion Capture, Multi-view Human Pose Estimation, Uncertainty Modeling, Synthetic Training
I Introduction
@@
The conventional 3D keypoint estimation process involves two stages. Firstly, 2D landmarks are extracted from each camera view. This is followed by triangulation using the known camera parameters to infer 3D points [2, 4]. However, the accuracy of such methods heavily relies on the precision of independent 2D predictions across views, which is problematic in scenarios with complex body-part interactions or severe occlusions. Outlier mitigation techniques such as RANdom-SAmple Consensus (RANSAC) [5] and refinement algorithms [1] offer some robustness but cannot fully address these inherent limitations. Recent advances in deep learning models that use cross-view fusion strategies [6, 7, 3] have yielded promising 3D pose estimation results. For instance, epipolar transformers [3] show the benefits of leveraging cross-view information using epipolar geometry. However, the scalability of such approaches is often hindered with additional cameras, requiring advanced techniques to maximize the agreement between several model outputs [7]. Another research direction aims to leverage rich temporal information [8, 6] to enhance pose estimation accuracy. For instance, numerous works show the impact of using large temporal context for monocular 3D pose estimation [9, 10]. Similarly, methods like MFT-Transformers [6], focusing on multi-view fusion and temporal modelings, can yield improvements over single-frame methods. However, such approaches require access to large annotated 3D datasets (multi-view video streams and corresponding 3D pose coordinates) during training, which is especially scarce in outdoor and in-the-wild settings. Furthermore, these models are often trained on limited pose and camera variations, hindering their generalization to novel views.
@@
To address the challenges of viewpoint scalability and reliance on 3D annotated training data, we introduce UPose3D, a new method for 3D human pose refinement. Our method leverages 2D keypoints and their uncertainties from two sources to improve robustness to outliers and noisy data. These sources include: a) direct 2D pose estimation from RGB images, and b) a pose compiler module that utilizes consistency across views and over time. Additionally, we introduce a new cross-view fusion strategy to ensure scalability to different numbers of cameras before our pose compiler. Specifically, we project the keypoints from all available views onto a reference view to obtain a 2D point cloud for each joint. These are then fed to a point cloud transformer module to learn cross-view representations. These features are then passed to a spatiotemporal encoder to efficiently process temporal and skeleton information from temporal windows. To train our pose compiler without relying on 3D annotated datasets, we generate synthetic data simulating realistic multi-view human pose recordings from a large-scale motion capture dataset. This approach promotes generalization across diverse camera configurations and postures, overcoming the limitations of real-world, multi-view 3D annotated datasets.
@@
To evaluate the performance of our proposed model, we use four widely used public datasets, Human3.6m [11], Rich [12], AMASS [13], and COCO-WholeBody [14, 15], across two separate experiment setups. More specifically, we assess the performance of our method in both in-distribution and out-of-distribution (OoD) settings to better evaluate generalizability to new environments and multi-view camera setups. These experiments demonstrate that our approach outperforms prior state-of-the-art solutions in OoD while achieving competitive performance in in-distribution settings. Next, we provide detailed ablation experiments demonstrating the impact of uncertainty modeling and our pose compiler in improving multi-view pose estimation robustness to outliers. Finally, we perform several experiments to showcase the impact of the number of camera views and larger time windows. In summary, our method achieves a high level of view-point flexibility and robustness without requiring direct 3D annotations.
Our contributions can be summarized as follows:
@@
    We present UPose3D, a 3D human pose estimation pipeline for multi-view setups that achieves state-of-the-art results in OoD settings and performs competitively in in-distribution evaluations.
@@
    Our method uses a novel uncertainty-aware 3D pose estimation algorithm that uses normalizing flows to leverage 2D pose distribution modeling. Experiments demonstrate that this approach is more effective than prior works using heatmaps.
@@
    We propose a new training strategy that relies only on synthetic multi-view motion sequences generated online from motion capture data, which allows our pipeline to scale to various camera and skeleton configurations.
II Related Works
3D Pose Estimation. Traditionally triangulation techniques like RANSAC have been used for 3D human pose estimation [2, 4]. However, these methods are generally not directly differentiable and their integration into deep learning pipelines is hindered. Therefore, recent research has explored more flexible, soft-predictive models, such as volumetric 3D keypoint representations [2] and cross-view feature fusion strategies [7]. Another approach [16] introduces a stochastic framework for human pose triangulation that relies on 3D pose hypothesis generation, scoring, and selection from 2D detection of several camera views. However, the accuracy of 2D pose detectors limits most such approaches. As a result, some methods incorporate epipolar geometry for pose consistency via self-supervised [17] and semi-supervised [18] learning. More advanced techniques have explored feature fusion strategies using epipolar lines across camera pairs [3, 4], showing significant improvements. However, these methods require complex re-projection in view pairs, limiting their scalability to any number of cameras or different camera placements due to the lack of camera variety in the training set.
To avoid the reliance on large number of parameters, a recent work [6] proposed a multi-view and temporal transformer network to perform an end-to-end 3D pose estimation without relying on camera parameters via leveraging cross-view feature fusion. Other works have conducted experiments on using human pose priors using generative models such as GANs [19] and Diffusion probabilistic models [20] to perform 3D pose estimations from a variety of inputs. However, most such approaches rely on 3D pose annotations from in-studio collected datasets to train their model.
Another research track explores the benefits of uncertainty modeling for 3D human pose estimation for enhancing the performance in occlusion-heavy scenarios [21, 22]. For example, Residual Log-likelihood Estimation (RLE) [23], has been proposed to model the underlying distribution of 2D keypoints in human pose estimation via regression. This approach leverages a re-parameterization technique and normalizing flows to learn keypoint uncertainty and achieves similar performance to heatmap-based [24, 25] and SimCC-based [26] techniques. One advantage of such uncertainty modeling techniques that we exploit in this paper is their ability to estimate the likelihood of keypoints in an image in a differentiable process, which other methods lack.
@@ II Related Works
III Method
Our method, Upose3D, determines the 3D coordinates of body joints from one or more consecutive image frames captured from multiple viewpoints. Figure 1 depicts an overview of our proposed method. At first, an RLE-based 2D pose estimator (Sec. III-A) extracts 2D keypoints and their corresponding uncertainties from the input images. The keypoints are then projected onto other views using epipolar geometry (Sec. III-B). From there, the pose compiler (Sec. III-C) refines the keypoint locations and their uncertainties by leveraging spatiotemporal and cross-view information. Lastly, using outputs from the 2D pose estimator and the pose compiler, an iterative refinement stage obtains the final 3D poses (Sec. III-D). We end this section by describing our multi-view training data synthesis (Sec. III-E).
III-A 2D Pose Estimation
The first step of our method is estimating 2D joint position in each image $\mathcal{I}_{i,t}$ from all camera views $i\in\mathcal{V}=\{1,2,\dots,V\}$ and frames $t\in\mathcal{T}=\{1,2,\dots,T\}$. Similar to [34, 35], our method implements a single layer Residual Log-likelihood Estimation (RLE) head [23] on top of an off-the-shelf backbone (e.g., CPN [36]). Aside from being computationally effective and robust to occlusion, RLE provides uncertainty $\hat{\sigma}$ for each joint position prediction $\hat{\mu}$. Specifically, the RLE predicts a distribution $P_{\Theta}(x|\mathcal{I})$ that models the probability of the ground-truth keypoints appearing in position $x$ using a normalizing flow model [37] with learnable parameters $\Theta$. The $\hat{\mu}$ produced by this first module is refined in the next stages by leveraging cross-view and temporal information. The estimated $\hat{\mu}$, $\hat{\sigma}$, along with the normalizing flow parameters $\Theta$, are also used during our final 3D keypoint estimation process in Sec. III-D.
III-B Cross-view Projection
@@ III-C Pose Compiler
Next, we concatenate the feature vectors from every frame and joint in each reference view $\{\textbf{f}_{i}\}$ with temporal and spatial (i.e., frame and skeleton joint ID) position embeddings. The results are then passed to our spatiotemporal encoder that implements an RLE head to provide position $\hat{\mu}^{\prime}_{i,k,t}$ and uncertainty $\hat{\sigma}^{\prime}_{i,k,t}$ estimations. Our spatiotemporal encoder is a transformer [41], with specific attention block modifications that accommodate the large dimensionality of the tensors containing all temporal and skeleton information. As detailed in Fig. 2, we replace the standard attention module with criss-cross attentions [33, 8], which approximates full temporal and spatial dependencies while being more memory efficient than full attention.
Accordingly, the cross-view input features $\{\textbf{f}_{i}\}$ are first projected into queries, keys, and values ($Q,K,V\in\mathbb{R}^{T\times J\times 2H}$) via a linear layer. Next, we divide them into temporal $Q_{T},K_{T},V_{T}\in\mathbb{R}^{T\times J\times H}$ and spatial groups $Q_{S},K_{S},V_{S}\in\mathbb{R}^{T\times J\times H}$. The temporal and spatial (skeleton joints) attentions are then calculated in two separate self-attention modules and concatenated before the next feed-forward layer and normalization. As a result of this operation, the receptive field of each transformer layer is the information residing on the spatial and temporal axis, and stacking multiple layers can approximate the full spatiotemporal attention without large computational overhead.
We train the RLE head of the spatiotemporal encoder using the same strategy as [23]. Finally, similarly to Sec. III-A, the RLE head simultaneously maximizes and learns a distribution $P_{\Theta^{\prime}}(x|\mathcal{C})$ that represents the probability of the appearance of a keypoint in position $x$ using a normalizing flow model with learnable parameters $\Theta^{\prime}$. In the next section, we use the estimated $\hat{\mu}^{\prime}$, $\hat{\sigma}^{\prime}$, and $\Theta^{\prime}$ for 3D pose estimation.
III-D 3D Pose Estimation
@@
III-E Multi-view Training and Data Synthesis
Our proposed pose compiler does not require 3D annotated data, given that it operates on point clouds. Therefore, we can use animation data to create a synthetic training set. In our In-Distribution (InD) experiments, we use the SMPL+H [43] body model and a Human3.6m [11] specific 17-joint regressor provided by [44]. For our OoD experiments, we use the SMPL-X [45] body model and train a whole-body 133-joint regressor from scratch synthetically on 1000 rendered multi-view images [46].
Let $r\in\mathbb{R}^{T\times 3}$, $\Theta\in\mathbb{R}^{T\times 55\times 3\times 3}$, and $\beta\in\mathbb{R}^{16}$ be root position, body joint rotation matrices, and shape parameters from human motion capture data. We begin our multi-view data generation by augmenting the shape parameters with Gaussian noise with a standard deviation equal to the standard deviation of all shape parameters within the datasets. Next, we apply mediolateral mirroring with a 50% chance and randomly rotate the motion sequence around its center. We pass the augmented $\{r,\Theta,\beta\}$ parameters to the forward-kinematic layer of the SMPL body model to obtain 3D vertices. Lastly, we use the dataset-specific joint regressor on the vertices to extract the 3D keypoints used in the next steps of our pipeline.
After obtaining the camera intrinsic and extrinsic parameters, we project the 3D body keypoints onto each camera view (see Fig. 3-b). We use these 2D keypoints as ground-truths $\mu_{g}$ to train our pose compiler (see Fig. 3-c). We then add 2D point corruptions to the 2D keypoints, including Gaussian noise with varying standard deviations, simulated occlusions with varying sizes and probabilities, mediolateral flipping, and occasional truncation effects (see Fig. 3-d). Next, we obtain the cross-view projected keypoints (see Fig. 3-e) via the algorithm described in Sec. III-B. Finally, we train our pose compiler using the ground-truth keypoints and point clouds containing noisy 2D data, as depicted in Fig. 4.
Figure 4: We illustrate the training routine of our pose compiler using synthetic data generated from human motions.
IV Experiments
IV-A Datasets
@@
We compare our proposed approach with prior works on the Human3.6m [11] dataset and measure how our method generalizes to different skeleton configurations and unseen outdoor environments using the COCO WholeBody [14] and RICH [12] datasets. The pose compiler is trained using motion capture data from the AMASS dataset [13] to simulate multi-view training data. Finally, we analyze the viewpoint scalability of our approach on the CMU Panoptic [47] dataset. The details of these datasets are as follows:
@@
Human3.6m. The Human3.6m [11] dataset is a standard benchmark for evaluating the performance of 3D human pose estimation solutions in multi-view [6] and single-view [9] settings. We report our InD performance using Protocol-I by training our 2D pose estimator on subjects 1, 5, 6, 7, and 8 and evaluating UPose3D on subjects 9 and 11. We use the same 3D ground-truth keypoints and camera parameters as [5].
@@
RICH. Real scenes, Interaction, Contact, and Humans (RICH) dataset [12] is a recently published dataset of multi-view videos with accurate markerless motion captured bodies and scenes. The dataset contains human bodies and 3D-scanned scenes obtained using markerless motion capture technology. The 3D annotations are provided in SMPL-x [45] format, from which we extract 3D ground-truth keypoints for evaluation. We report our OoD performance on the test set, which contains 7 subjects in 52 scenarios and 3 environments, including a construction site, a gym, and a lecture hall.
COCO WholeBody. The COCO WholeBody dataset is a large-scale whole-body pose estimation dataset with over 250K samples [14]. This dataset is an extension of Common Objects in COntext (COCO) [15] dataset with the same training and testing splits. The dataset provides 133 2D keypoints (17 for body, 6 for feet, 68 for face, and 42 for hands) on in-the-wild images. We use this dataset to train our 2D pose estimator during OoD experiments.
@@
CMU Panoptic. The CMU Panoptic [47] dataset is a large collection of human body poses and interactions recorded from multiple views. The dataset provides videos from 480 VGA and 30 HD cameras and ground-truth 3D poses obtained from markerless 2D pose estimation and triangulation. We use a small subset of the validation set in line with [2] to analyze the scalability of our approach with the addition of camera views.
@@
AMASS. The Archive of Motion Capture as Surface Shapes (AMASS) [13] is a large collection of 3D human motion capture datasets. It contains over 40 hours of recording from 300 subjects spanning 11,000 actions. An important contribution of this dataset is the unification of 3D human bodies under the SMPL [43] and SMPL-x [45] parametric models used to train strong motion and pose priors such as VPoser [45] and HuMoR [48]. We only use the training set of this dataset following prior works [48, 45].
IV-B Evaluation Metrics

### v3

@@ Abstract
We introduce UPose3D, a novel approach for multi-view 3D human pose estimation, addressing challenges in accuracy and scalability. Our method advances existing pose estimation frameworks by improving robustness and flexibility without requiring direct 3D annotations. At the core of our method, a pose compiler module refines predictions from a 2D keypoints estimator that operates on a single image by leveraging temporal and cross-view information. Our novel cross-view fusion strategy is scalable to any number of cameras, while our synthetic data generation strategy ensures generalization across diverse actors, scenes, and viewpoints. Finally, UPose3D leverages the prediction uncertainty of both the 2D keypoint estimator and the pose compiler module. This provides robustness to outliers and noisy data, resulting in state-of-the-art performance in out-of-distribution settings. In addition, for in-distribution settings, UPose3D yields performance rivalling methods that rely on 3D annotated data while being the state-of-the-art among methods relying only on 2D supervision.
Keywords:
Markerless Motion Capture Multi-view Human Pose Estimation Uncertainty Modeling Temporal Learning
1 Introduction
@@
The conventional 3D keypoint estimation process involves two stages. Firstly, 2D landmarks are extracted from each camera view. This is followed by triangulation using the known camera parameters to infer 3D points [19, 66]. However, the accuracy of such methods heavily relies on the precision of independent 2D predictions across views, which is problematic in scenarios with complex body-part interactions or severe occlusions. Outlier mitigation techniques such as RANdom-SAmple Consensus (RANSAC) [43] and refinement algorithms [60] offer some robustness but cannot fully address these inherent limitations. Recent advances in deep learning models that use cross-view fusion strategies [55, 47, 14] have yielded promising 3D pose estimation results. For instance, Epipolar Transformers [14] shows the benefits of leveraging cross-view information using epipolar geometry. However, the scalability of such approaches is often hindered with additional cameras, requiring advanced techniques to maximize the agreement between several model outputs [47]. Another research direction aims to leverage rich temporal information [58, 55] to enhance pose estimation accuracy. For instance, numerous works show the impact of using large temporal context for monocular 3D pose estimation [70, 48]. Similarly, methods like MFT-Transformers [55], focusing on multi-view fusion and temporal modelings, can yield improvements over single-frame methods. However, such approaches require access to large annotated 3D datasets (multi-view video streams and corresponding 3D pose coordinates) during training, which is especially scarce in outdoor and in-the-wild settings. Furthermore, these models are often trained on limited pose and camera variations, hindering their generalization to novel views.
@@
To address the challenges of viewpoint scalability and reliance on 3D annotated training data, we introduce UPose3D, a new method for 3D human pose estimation. Our method leverages 2D keypoints and their uncertainties from two sources to improve robustness to outliers and noisy data. These sources include: a) direct 2D pose estimation from RGB images, and b) a pose compiler module that utilizes consistency across views and over time. Additionally, we introduce a new cross-view fusion strategy to ensure scalability to different numbers of cameras before our pose compiler. Specifically, we project the keypoints from all available views onto a reference view to obtain a 2D point cloud for each joint. These are then fed to a point cloud transformer module to learn cross-view representations. These features are then passed to a spatiotemporal encoder to efficiently process temporal and skeleton information from temporal windows. To train our pose compiler without relying on 3D annotated datasets, we generate synthetic data simulating realistic multi-view human pose recordings from a large-scale motion capture dataset. This approach promotes generalization across diverse camera configurations and postures, overcoming the limitations of real-world, multi-view 3D annotated datasets.
@@
To evaluate the performance of our proposed model, we use four widely used public datasets, Human3.6m [18], RICH [16], HUMBI [64, 65], and CMU Panoptic [22], across two separate experiment setups. More specifically, we assess the performance of our method in both in-distribution (InD) and out-of-distribution (OoD) settings to better evaluate generalizability to new environments and multi-view camera setups. These experiments demonstrate that our approach outperforms prior state-of-the-art solutions in OoD while achieving competitive performance in InD settings. Next, we provide detailed ablation experiments demonstrating the impact of uncertainty modeling and our pose compiler in improving multi-view pose estimation robustness to outliers. Finally, we perform several experiments to showcase the impact of the number of camera views and larger time windows. In summary, our method achieves a high level of view-point flexibility and robustness without requiring direct 3D annotations. Our contributions can be summarized as follows:
@@
    We present UPose3D, a 3D human pose estimation pipeline for multi-view setups that achieves state-of-the-art results in OoD settings and performs competitively in InD evaluations.
@@
    Our method uses a novel uncertainty-aware 3D pose estimation algorithm that uses normalizing flows to leverage 2D pose distribution modeling. Experiments demonstrate that this approach is more effective than prior works using heatmaps in terms of accuracy and computational costs. This also allows our pipeline to scale to different camera setups with constant runtime.
@@
    We propose a new training strategy that relies only on synthetic multi-view motion sequences generated online from motion capture data.
2 Related Works
3D Pose Estimation. Traditionally, triangulation techniques like RANSAC have been used for 3D human pose estimation [19, 66]. However, these methods are generally not directly differentiable and their integration into deep learning pipelines is hindered. Therefore, recent research has explored more flexible, soft-predictive models, such as volumetric 3D keypoint representations [19] and cross-view feature fusion strategies [47]. Another approach [2] introduces a stochastic framework for human pose triangulation that relies on 3D pose hypothesis generation, scoring, and selection from 2D detection of several camera views. However, the accuracy of 2D pose detectors limits most such approaches. As a result, some methods incorporate epipolar geometry for pose consistency via self-supervised [27] and semi-supervised [63] learning. More advanced techniques have explored feature fusion strategies using epipolar lines across camera pairs [14, 66], showing significant improvements. However, these methods require complex re-projection in view pairs, limiting their scalability to a large number of cameras or with different placements due to lack of variety in the training set.
To avoid relying on large parameters, a recent work [55] proposed a multi-view temporal transformer network for end-to-end 3D pose estimation by leveraging cross-view feature fusion. Other works have conducted experiments on using human pose priors using generative models such as GANs [25] and Diffusion probabilistic models [15] to perform 3D pose estimations from a variety of inputs. However, most such approaches rely on 3D pose annotations from in-studio collected datasets to train their model. Another research track explores the benefits of uncertainty modeling for 3D human pose estimation for enhancing the performance in occlusion-heavy scenarios [3, 29]. For example, Residual Log-likelihood Estimation (RLE) [31], has been proposed to model the underlying distribution of 2D keypoints in human pose estimation via regression. This approach leverages a re-parameterization technique and normalizing flows to learn keypoint uncertainty and achieves similar performance to heatmap-based [57, 4] and SimCC-based [33] techniques. We exploit the ability of RLE to estimate a differentiable likelihood of keypoints, which other methods lack.
@@ 2 Related Works
3 Method
Upose3D determines the 3D coordinates of body joints from one or more consecutive multi-view frames. Figure 1 depicts an overview of our proposed method. At first, an RLE-based 2D pose estimator (Sec. 3.1) extracts 2D keypoints and their corresponding uncertainties from the input images. The keypoints are then projected onto other views using epipolar geometry (Sec. 3.2). From there, the pose compiler (Sec. 3.3) refines the keypoint locations and their uncertainties by leveraging spatiotemporal and cross-view information. Lastly, using outputs from the 2D pose estimator and the pose compiler, an iterative algorithm obtains the final 3D poses (Sec. 3.4). We end this section by describing our multi-view training data synthesis (Sec. 3.5).
3.1 2D Pose Estimation
The first step of our method is estimating 2D joint position in each image $\mathcal{I}_{i,t}$ from all camera views $i\in\mathcal{V}=\{1,2,\dots,V\}$ and frames $t\in\mathcal{T}=\{1,2,\dots,T\}$. Similar to [30, 8], our method implements a single layer Residual Log-likelihood Estimation (RLE) head [31] on top of an off-the-shelf backbone (e.g., CPN [5]). Aside from being computationally effective and robust to occlusion, RLE provides uncertainty $\hat{\sigma}$ for each joint position prediction $\hat{\mu}$. Specifically, the RLE predicts a distribution $P_{\Theta}(x|\mathcal{I})$ that models the probability of the ground truth appearing in position $x$ using a normalizing flow model [52] with learnable parameters $\Theta$. The $\hat{\mu}$ produced by this module is refined in the next stages by leveraging cross-view and temporal information. The estimated $\hat{\mu}$, $\hat{\sigma}$, along with the normalizing flow parameters $\Theta$, are used for 3D keypoint estimation in Sec. 3.4.
3.2 Cross-view Projection
@@ 3.3 Pose Compiler
Next, we concatenate the feature vectors from every frame and joint in each reference view $\{\textbf{f}_{i}\}$ with temporal and spatial position embeddings, i.e., frame and skeleton joint ID. The results are then passed to our spatiotemporal encoder that implements an RLE head to provide position $\hat{\mu}^{\prime}_{i,k,t}$ and uncertainty $\hat{\sigma}^{\prime}_{i,k,t}$ estimations. Our spatiotemporal encoder is a transformer [61], with specific attention block modifications that accommodate the large dimensionality of the tensors containing all temporal and skeleton information. As detailed in Fig. 2, we replace the standard attention module with criss-cross attentions [17, 58], which approximates full temporal and spatial dependencies while being more memory efficient than full attention. We train the RLE head of the spatiotemporal encoder using the same strategy as [31]. Finally, similarly to Sec. 3.1, the RLE head simultaneously maximizes and learns a distribution $P_{\Theta^{\prime}}(x|\mathcal{C})$ that represents the probability of the appearance of a keypoint in position $x$ using a normalizing flow model with learnable parameters $\Theta^{\prime}$. In the next section, we use the estimated $\hat{\mu}^{\prime}$, $\hat{\sigma}^{\prime}$, and $\Theta^{\prime}$ for 3D pose estimation.
3.4 3D Pose Estimation
@@
3.5 Training and Multi-view Data Synthesis
Our proposed pose compiler does not require 3D annotated data, given that it operates on point clouds. Therefore, we can use any animation data to create a synthetic training set. The synthesis of multi-view training data starts by randomly positioning several cameras in 3D space to look roughly at the human body’s center. During this process, we limit the camera height, distance, tilt, pitch, and yaw within a representative range of a normal multi-view recording setup. Next, we extract the 3D landmarks and project them onto each camera to obtain ground-truth 2D keypoints $\mu_{g}$. We then corrupt these keypoints through extensive augmentations. Refer to our Supplementary Materials for more details on our data synthesis. Finally, we obtain the point clouds training data from the corrupted 2D keypoint using the cross-view projection from Sec. 3.2.
4 Experiments
4.1 Datasets
@@
We compare our proposed approach with prior works on the Human3.6m [18] dataset and measure how our method generalizes to different skeleton configurations and unseen outdoor environments on the RICH [16] dataset. We use the COCO WholeBody [21, 36] and MPII [1] for training our 2D pose estimator in different scenarios (refer to the Supplementary Materials for more information [69]). The pose compiler is trained using motion capture data from the AMASS dataset [42] to simulate multi-view training data. Finally, we analyze the viewpoint scalability of our approach on the CMU Panoptic [22] and HUMBI [64, 65] datasets. The details of these datasets are as follows.
@@
Human3.6m. The Human3.6m [18] dataset is a standard benchmark for evaluating the performance of 3D human pose estimation solutions in multi-view [55] and single-view [70] settings. We report our InD performance in Protocol-I using subjects 1, 5, 6, 7, and 8 for training and 9 and 11 for testing.
@@
RICH. Real scenes, Interaction, Contact, and Humans (RICH) dataset [16] is a dataset of multi-view videos with accurate bodies and 3D-scanned scenes obtained using markerless motion capture technology. We extract 3D ground-truth keypoints from the provided SMPL-x [46] bodies for evaluation. We report OoD performance on the test set, which contains 7 subjects in 52 scenarios and 3 environments, including a construction site, a gym, and a lecture hall.
@@
CMU Panoptic. The CMU Panoptic [22] dataset is a large collection of human body poses and interactions recorded from multiple views. The dataset provides videos from 480 VGA and 30 HD cameras and ground-truth 3D poses obtained from markerless 2D pose estimation and triangulation. We use a small subset of the validation set in line with [19] to analyze UPose3D’s camera scalability.
HUMBI. The HUman Multiview Behavioral Imaging (HUMBI) [64, 65] dataset is a large-scale multi-view dataset designed for modeling and reverse-rendering of the human body in different expressions. It contains 107 HD recordings of 772 subjects with natural clothing. We use this dataset only for OoD evaluations during our scalability analysis. Specifically, we evaluate on the first 80 subjects by extracting ground-truth 3D keypoints from the provided SMPL [39] parameters.
@@
AMASS. The Archive of Motion Capture as Surface Shapes (AMASS) [42] is a large collection of 3D human motion capture datasets. It contains over 40 hours of recording from 300 subjects spanning 11,000 actions. We only use the training set of this dataset following prior works [50, 46, 7]. Since the extraction of landmarks requires a specialized joint regressor for each skeleton configuration, in our InD experiments, we use the SMPL+H [39] body model and a 17-joint regressor provided by [44]. For our OoD experiments, we use the SMPL-x [46] body model and a whole-body 133-joint regressor from [7].
4.2 Evaluation Metrics